OpenStudy (anonymous):

How are these functions inverses: f(n)=2(n-2)^3 and g(n)=4+(4n^1/3)/2 (or cubic root: 4n)

OpenStudy (amistre64):

if f(g(n)) = g(f(n)) = n they is inverses

OpenStudy (anonymous):

\[f(n)=2(n-2)^{3} \] \[g(n)= (4+\sqrt[3]{4n})\div2 \]

OpenStudy (anonymous):

but can you show me how?

OpenStudy (anonymous):

I've tried matching f(n) to g(n) but I just can't D: eek

OpenStudy (anonymous):

and I don't know how to for this type

OpenStudy (amistre64):

plug g(n) into f(n) and calulate it out; if it reduces to "n" its an inverse in that direction

OpenStudy (anonymous):

how about using this method: y=2(n-2)^3 n=2(y-2)^3 .....

OpenStudy (amistre64):

that method might be useful if they both equate in the end \[f(n)=2(n-2)^{3}\] \[\frac{f(n)}{2}=(n-2)^{3}\] \[\sqrt[3]{\frac{f(n)}{2}}=n-2\] \[\sqrt[3]{\frac{f(n)}{2}}+2=n\]

OpenStudy (anonymous):

but don't they equate since they're both inverses of each other?

OpenStudy (anonymous):

I keep getting: \[y=(\sqrt[3]{n+8})\div2\]

OpenStudy (amistre64):

f(n)=2(n-2)^3 g(n)=4+(4n^1/3)/2 f(g(n)) = 2((4+(4n^1/3)/2)-2)^3 f(g(n)) = 2((2+(4n^1/3)/2))^3 f(g(n)) = 2((2+(4n^1/3)/2))^3 1 3 3 1 8 4 2 1 1 4'1/3n 4'2/3n^2 4'3/3n^3 8 12 im going blind with this thing

OpenStudy (amistre64):

who says they are inverses of each other?

OpenStudy (anonymous):

on my worksheet it says these two are inverses of each other: http://www.kutasoftware.com/FreeWorksheets/Alg2Worksheets/Function%20Inverses.pdf

OpenStudy (anonymous):

#6

OpenStudy (amistre64):

"State if the given functions are inverses." that does not say that they are indeed inverses

OpenStudy (anonymous):

oh the answers are at the bottom

OpenStudy (amistre64):

.... :/ thnx lol

OpenStudy (amistre64):

it might be easier to test g(f(n)) to start with

OpenStudy (anonymous):

ohh, okay

OpenStudy (anonymous):

so I have: \[f(g(n)) = 2((4+\sqrt[3]{4n}\div2)-2)\]

OpenStudy (amistre64):

f(n)=2(n-2)^3 g(n)=(4+(cbrt(4n))/2 g(f(n)) = (4+(cbrt(4(2(n-2)^3)))/2 g(f(n)) = (4+(cbrt(8 (n-2)^3))/2 g(f(n)) = (4+2cbrt((n-2)^3)/2 g(f(n)) = (4+2(n-2))/2 g(f(n)) = 2+(n-2) = n

OpenStudy (anonymous):

^3

OpenStudy (anonymous):

where did you get the 2 from: g(f(n)) = (4+2cbrt((n-2)^3)/2

OpenStudy (anonymous):

ohh, nevermind: 8^1/3

OpenStudy (amistre64):

trying to determine the other route is a pain

OpenStudy (anonymous):

I'm sorry D: Question: What happened to the cubic root here: g(f(n)) = (4+2(n-2))/2

OpenStudy (amistre64):

cbrt..^3 cancel each outher out f(-1) = 2(-1-2)^3 = -54 g(-54)=(4+cbrt(4(-54)))/2 does not equal -1

OpenStudy (anonymous):

ohhh

OpenStudy (amistre64):

i believe you have a restricted domain inverse with these

OpenStudy (amistre64):

f(5) = 54 g(54) = 5 f(-1) = -54 g(-54) not= -1

OpenStudy (amistre64):

does the "n" signify natural numbers?

OpenStudy (anonymous):

ohh... so I have to use restrictions? And last question: Why is it: g(f(n)) = 2+(n-2) = n In the end I got: 4+2(n-2)/2=6(n-2)/2...

OpenStudy (anonymous):

I don't know because they don't state anything about what 'n' means... I just thought it's just a variable like x

OpenStudy (amistre64):

x is traditioanlly used to define any real variable; and n is used to define any natural number

OpenStudy (anonymous):

ohh.. so what happens? o.O

OpenStudy (amistre64):

g(f(n)), for n>=2, these are inverses

OpenStudy (anonymous):

so I use restrictions? How did you know it would be n>=2? Is it because of the (n-2)?

OpenStudy (amistre64):

\[g(f(n)) = 2+\frac 12 \sqrt[3]{4(2(n-2)^3)}\] \[g(f(2)) = 2+\frac 12 \sqrt[3]{4(2(2-2)^3)}=2+\frac 12 \sqrt[3]{0}=2\] \[g(f(3)) = 2+\frac 12 \sqrt[3]{4(2(3-2)^3)}=2+\frac 12 \sqrt[3]{8}=2+1=3\] \[g(f(1)) = 2+\frac 12 \sqrt[3]{4(2(1-2)^3)}\]\[g(f(1)) = 2+\frac 12 \sqrt[3]{4(2(-1)^3)}\]\[g(f(1)) = 2+\frac 12 \sqrt[3]{-8}\]\[g(f(1)) = 2+\frac 12 (-2)=1\] harumpf!! works for 1 what about 0? \[g(f(0)) = 2+\frac 12 \sqrt[3]{4(2(0-2)^3)}\]\[g(f(0)) = 2+\frac 12 \sqrt[3]{4(2(-8))}\]\[g(f(0)) = 2+\frac 12 \sqrt[3]{8(-8))}\]\[g(f(0)) = 2+\frac 12 (2)(-2)\]\[g(f(0)) = 2-2=0\] -1 perhaps? \[g(f(-1)) = 2+\frac 12 \sqrt[3]{4(2(-1-2)^3)}\]\[g(f(-1)) = 2+\frac 12 \sqrt[3]{8(-27)}\]\[g(f(-1)) = 2+\frac 12 (2)(-3)\]\[g(f(-1)) = 2-3=-1\] well, it appears f(n) is at least the inverse of g(n)

OpenStudy (amistre64):

do the same setup for f(g(n)) and see if you get "n" back

OpenStudy (anonymous):

ohh by trial and error of substituting the 'n' with a number?

OpenStudy (anonymous):

btw, thank you so much!

OpenStudy (amistre64):

i was testing out some stuff, the n>=2 might be wrong; but try it out for f(g(n)) to see if it applies